Notes on the Confluence Property of

Term Rewriting Systems and the λ-calculus

Computation Structures Group Memo 321

Zena M. Ariola

This report describes research done at the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for this work has been provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-89-J-1988 (MIT) and N0039-88-C-0163 (Harvard).